OpenStudy (anonymous):

Differntiate A=A(1/2)^5

OpenStudy (anonymous):

\[A(1/2)^5\]

OpenStudy (agent0smith):

\[\huge A= A_o \left( \frac{ 1 }{2 } \right) ^ {t/20} \] i think you need to differentiate from here. Then plug in your t value.

OpenStudy (anonymous):

Ohhhh I c

OpenStudy (anonymous):

ill try that

OpenStudy (anonymous):

0.1617....

OpenStudy (agent0smith):

for Ao you can plug in the (Ao/32) since there's only 1/32 remaining.

OpenStudy (anonymous):

Is it 5(1/2)^99/20

OpenStudy (agent0smith):

How'd you get that? That's not the derivative.

OpenStudy (anonymous):

well I took the exponent and put it in front and subtracted 1 form the exponent

OpenStudy (anonymous):

isnt that the rule?

OpenStudy (agent0smith):

No... you don't have a simple power function. you have one of these exponential a^x functions http://tutorial.math.lamar.edu/Classes/CalcI/DiffExpLogFcns_files/eq0031MP.gif

OpenStudy (anonymous):

A((1/2)^t/20)(ln)(1/2)

OpenStudy (anonymous):

\[A((t/20)^{t/20}\ln(1/2) \]

OpenStudy (anonymous):

0.0224.....

OpenStudy (agent0smith):

I got an answer much smaller than that

OpenStudy (agent0smith):

It's not quite as simple as just doing the rule with a^x

OpenStudy (anonymous):

i think im just gonna save this question for my math friend

OpenStudy (anonymous):

thx anyways

OpenStudy (agent0smith):

Well the answer should be \[\Large \frac{ dA }{ dt } = -3.38 \times 10 ^{-5} A_o\] but it takes a bit of work to get to.